OpenStudy (anonymous):

What is the center and radius of the following circle: (x – 5)2 + (y + 2)2 = 16

OpenStudy (anonymous):

I'll define u and v the following way: \[u = x-5; \; v = y+2\] Now I can say that \[u² + v² = r²\]and this equation describes a circle with radius r where r = 4 and the center is at (u=0, v=0). In terms of x and y, we can describe the location of the center as (5,-2).